OpenStudy (anonymous):

Simplify the radical expression. - square root of 32

OpenStudy (wearehiding):

\(4 sqrt(2)\)

OpenStudy (mathstudent55):

What is the largest perfect square factor of 32?

OpenStudy (jdoe0001):

\(\bf {\color{brown}{ 32\to 2\cdot 2\cdot 2\cdot 2\cdot 2\to 2^2\cdot 2^2\cdot 2}}\qquad thus \\ \quad \\ -\sqrt{32}\implies -\sqrt{2^2\cdot 2^2\cdot 2}\implies ?\)

OpenStudy (mathmate):

@WeAreHiding Please read: http://openstudy.com/code-of-conduct